Circular No. 5691

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1987A IN THE LARGE MAGELLANIC CLOUD A. P. S. Crotts, Columbia University; and W. E. Kunkel, Carnegie

Observatories, write: “The light echo from the circumstellar nebula, discovered in 1989 March extending some 2′′ from SN 1987A (IAUC 4791), has disappeared sometime between 1992 Jan. 15 and Mar. 20 UT (consistent with Crotts 1990, Proc. ESO/EIPC Workshop on “Supernova 1987A and Other Supernovae”, eds. Danziger and Kjar, p. 559; and somewhat at variance with several other models). The echo was seen clearly as a bridge extending approximately from star 2 to star 3 and passing just south of the supernova on 1991 Jan. 26. By 1992 Jan. 15, it had faded considerably to a barely perceptible patch in the vicinity of the supernova, and by 1992 Mar. 20 was no longer visible. This disappearance is below the level of 2 × 10^{−15} erg s^{−1} cm^{−2} A^{−1} arcsec^{−2}, as determined in the continuum at 612.0 nm. It has not reappeared in observations at several epochs since then. This behavior was similar in several other continuum bands, and corresponds to a maximal extent of the nebula behind the supernova of ∼ 2.2 lyr. The nebulosity is still visible in recombination lines, although dimming. This disappearance allows the photometry of SN 1987A to be determined more accurately in the continuum, where it is very faint, corresponding to V = 19.9, with r.m.s. errors of ∼ 0.2 mag, on 1992 Nov. 19. This sets limits on the presence of a stellar source, such as a companion star, corresponding to ∼ 2.5 M_{⊙} on the main sequence. The disappearance of the circumstellar echo allows precise subtraction of the flux from the remaining stars. This will allow an accurate determination of the evolution of the echo, hence the three-dimensional geometry of the bipolar nebula. We are preparing these data for publication.”

PERIODIC COMET SCHAUMASSE (1992x) Total visual magnitude estimates: 1992 Nov. 24.80 UT, 14.3 (A. Naka-

mura, Kuma, Japan, 0.60-m reflector); 17.26, 12.7 (A. Hale, Las Cruces, NM, 0.41-m reflector); 24.88, 12.2 (F. Van Loo, Mont Ventoux, France, 0.20-m reflector); 27.07, 12.0 (J. E. Bortle, Stormville, NY, 0.50-m reflector); 29.84, 12.5 (T. Vanmunster, Landen, Belgium, 0.35-m reflector); 1993 Jan. 12.12, 10.7 (C. Spratt, Victoria, BC, 0.20-m reflector).

1993 January 15 Daniel W. E. Green